1. If a new building has a replacement cost of $750,000, what is the minimum amount of insurance required to comply with an 80% coinsurance requirement?

Questions 2-5 apply to the following situation:

A commercial building that has a value of $600,000 is insured on a Broad Perils policy with a coverage limit of $400,000. and a deductible of $1,000.

2. There is a fire loss in the amount of $250,000. There is an 80% co-insurance clause. How much will the insurance company pay?

3. During a very cold part of an Erie winter, ice forms under the shingles on the roof of this building. During a thaw, this ice melts and backs up under the shingles, causing the melting water to enter the structure and to damage the interior ceiling of this building. The amount of this melting water damage was $40,000. Would this policy respond to this loss? If so, how much would be paid?

4. There is a claim for wind damage in the amount of $2,000. How much will the company pay?

5. An airplane crashes into the building and totally destroys it. How much will the insurance company pay for this total loss?

Pro-Rate Problem

6. There is a Basic Perils insurance policy in the amount on an unprotected restaurant. The building has a value of $1,000,000. This risk is divided among three companies. There is a deductible, but that is not part of this problem. Company A carries $450,000 of the risk, while Company B carries $350,000, and Company C carries a limit of $100,000. There is an 80% co-insurance clause. How much will each company pay for a $500,000 fire loss?
Answer #1

1.$ 600000(750000*0.8)

2.$200000(250000*0.8)

3.Yes, This policy should respond to this loss. Claim = 40000*0.667=$26680

4.$1334(2000*0.667)

5.$399000

6.Net Claim = 500000*0.8=$400000

A's Share=400000*45/90=$200000

B's Share= 400000*35/90=$155555.56

C's Share=400000*10/90=$44444.44